OpenStudy (anonymous):

How many sides does a polygon have if the sum of its angle measures is 2700?

OpenStudy (saifoo.khan):

Use the interior angle formula.

OpenStudy (anonymous):

gr8 like this only.......kp it up

OpenStudy (saifoo.khan):

\[2700 = \frac{(n-2)\times 180}{n}\]Solve for n.

OpenStudy (anonymous):

ok thank you so much!!!!

OpenStudy (saifoo.khan):

You welcome. please hit good answer.

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

ty.

OpenStudy (anonymous):

no problem

OpenStudy (saifoo.khan):

^_^

OpenStudy (anonymous):

:)

OpenStudy (zarkon):

why do you have an n in the denominator of your formula?

OpenStudy (anonymous):

i have 15n over 15n-30 which = 1. whats the next step?

OpenStudy (zarkon):

solve \[2700 = (n-2)\times 180\] for n

OpenStudy (anonymous):

ok thank you.

OpenStudy (anonymous):

i got n=17

OpenStudy (anonymous):

which means it has 17 sides right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thank you very much. that was easier for me

OpenStudy (zarkon):

good